Stock Values (LO1) The Stopperside Wardrobe Co. just paid a dividend of $1.45 per share on...

  1. Stock Values (LO1) The Stopperside Wardrobe Co. just paid a dividend of $1.45 per share on its stock. The dividends are expected to grow at a constant rate of 6 percent per year indefinitely. If investors require an 11 percent return on The Stopperside Wardrobe Co. stock, what is the current price? What will the price be in three years? In 15 years?
Answer #1

Current price=D1/(Required return-Growth rate)

=(1.45*1.06)/(0.11-0.06)

=$30.74

P3=Current price*(1+Growth rate)^3

=30.74*(1.06)^3

=$36.61(Approx).

P15=Current price*(1+Growth rate)^15

=30.74*(1.06)^15

=$73.67(Approx).